3. A box contains 4 bad and 6 good tubes. Two are drawn out from the box at a time. One is tested and found to be good. What is the probability that the other one is also good?


1
Expert's answer
2022-02-23T12:52:28-0500

After the first good tube picked up,

number of good tubes left = 5

number of bad tubes left = 4

The probability that the second tube picked up is good is


"\dfrac{5}{5+4}=\dfrac{5}{9}"